OpenStudy (anonymous):

Find two numbers whose diffrence is 22, and the larger number is 2 less then twice the smaller number.

OpenStudy (anonymous):

x-y = 22 x= 2y-2 2y-2-y=22 y=24 back into first equation x-24=22 x=46